A forensic psycho! gst tests the e tent to which the age o a crim nal predicts the age of the victim Y for nori lent crimes. The psychol st uses the case files to record the age of five ฌ minals and the age of the tim n t se cases. The hypothetical data are isted in the Age of Criminal Age of Victm 21 25 28 17 17 (a) Compute the method of least squares to find the equation of the regression line. (Round your answers to three dedimal places.) (b) Use the regression equation to determine the predicted age of a victim of a nonviolent crime when the criminal is 20 years old. (Round your answer to three decimal places. ears old Additional Materials -2 peints Pvters5tet3 16.EU21. A instructor measured quz scores and he number of hours studying among ฮ sample af 2u colege students if SSAY 42 SSx 91 My 7 and Mx 4, he mat is he regression cquation or this samp e? Round your num enc hrec decimal places. values Addilional Malerials -1 peints Prvder&Stets3 16 E 023 A researcher tested whethar time of day could predict mond in a sample of 15 college students. If 55ani208, then what is the standard eror of estimate in this sample?
Answer #1

17)

(x-x) 100.0000 1.0000 25.0000 36.0000 100.0000 262.0000 (y-y) 4.0000 1.0000 16.0000 0.0000 25.0000 46.0000 (x-x)(y-y 20.0000 1.0000 20.0000 0.0000 50.0000 89.0000 S.No 21 26 2 28 17 13 115 23.0000 4 17 110 22.0000 Total Mean SSY sample size here 22.0000 23.0000 from above 262.0000 Σ(X-X) (Y-Y)= 89.0000 slope- 0.3397 intercept- 14.1870 hence line equation: 14.187 0.340 x

from above:

Yhat=0.340x+14.187

b)predicted value =0.340*20+14.187=20.987

18)

here slope b1=42/94=0.447

and intercept bo=7-0.447*4=5.212

therefore line equation Y =0.447x+5.212 ( please try y=0.447x+5.213 if this comes wrong)

19)

std error =sqrt(SSE/(n-2))=sqrt(208/13)=4